A survey conducted about job satisfaction showed that 20% of workers are not happy with their current jobs. Two workers are selected at random, and it is observed whether or not they are happy with their current jobs. What is the probability that of these two workers, at least one of them is happy with their current job?
Answer by greenestamps(13200)



The opposite of "at least one of the two is happy" is "both are unhappy".

The probability that both are unhappy is (0.2)(0.2) = 0.04; so the probability that at least one is happy is 1-0.04 = 0.96.

An alternative method for calculating the probability....

P(both happy) = (0.8)(0.8) = 0.64
P(A happy, B not happy) = (0.8)(0.2) = 0.16
P(A not happy, B happy) = (0.2)(0.8) = 0.16

P(at least one happy) = 0.64+0.16+0.16 = 0.96